# Toward the Categorical Data Map

### v3

@@ III-B Projecting Categorical Data
The categorical data map should allow cluster similar categorical subsets and separating outliers addressing (R1) and (R2). Thus, we rely on dimensionality reduction (DR) to create a scatterplot-like visualization.
@@
Prepocessing: We convert all data items into a set representing their categorical data values. We define the set of all attributes as $\mathcal{A}\coloneqq\{a_{1},a_{2},\dots,a_{|\mathcal{A}|}\}$ and the possible categories associated with attribute $a_{i}$ as the set $\mathcal{C}_{i}\coloneqq\{c_{i}^{1},c_{i}^{2},\dots,c_{i}^{j}\}$ with $i,j\in\mathbb{N}$. Thus, $\left|\mathcal{A}\right|$ is the cardinality of a set representing a data item, i.e., the number of attributes since a data item has one category associated with each attribute. From a practical point of view, we make sure that all categories have a unique descriptor across all attributes. Thus, $\Sigma_{a_{i}\in\mathcal{A}}\,|\mathcal{C}_{i}|$ is the number of unique categories across all dimensions and the size of our vector representation for one-hot encoding (see subsection III-A). With this, we can use the sets $A$ and $B$, each describing the categories of a data item to define similarity. The three set-based distance measures described in subsection III-A can be applied directly. We also include one-hot encoding, converting each categorical value to a new binary dimension and using classical distance measures, such as Euclidean or Manhattan distance, to derive a dissimilarity relationship. We can use a combination of a data representation and a matching distance measure to calculate a distance matrix used with DR methods.
Subset-Based Representation of Categorical Data: Since categorical datasets can contain many duplicates, projecting each data point individually, using a DR method would lead to multiple data points being projected to the same position. The main reason is that the distance of identical points is zero. For iterative loss-based projection methods, this distance is rarely optimized to zero, which introduces distortions for this type of setup. Other methods plot identical points in the same location, resulting in overplotting. A second reason is that we want to show the subsets represented by a point; thus, showing all points individually is not our goal. Reducing the number of data points also improves the runtime of projection algorithms. The remaining data item describes the prototype of the represented data subset. Thus, rather than projecting the individual data points, we remove all duplicates and project one data point for each existing combination of attribute values, i.e., for each categorical subset. The size of this subset is later represented visually (see subsection III-C).
@@
Projection Methods: DR techniques are a set of non-/linear transformation methods with which a dataset’s dimensionality can be reduced. We selected the following DR methods:
Multidimensional Scaling (MDS): This method describes a set of linear and nonlinear DR techniques. They preserve pairwise distances. Multiple criteria are possible; Kruskal’s stress optimization criterion is usually used [59].
t-distributed Stochastic Neighbor Embedding (t-SNE): This technique is a nonlinear DR method. It enforces cluster structure by the divergence of a probability distribution of pairs of high-dimensional objects to keep similar data items near and dissimilar ones apart [45].
@@
Multiple Correspondence Analysis (MCA): This method is the categorical equivalent of PCA. It reduces categorical data to numeric values by quantifying the data items and their categories. MCA creates groups of items that are similar according to their categories. Objects sharing the same categories are placed close together, and objects with differing categories placed are far apart [25].
For the Categorical Data Map, we selected MDS and t-SNE. Both methods create a projection using a distance matrix. We chose the two methods based on their popularity and common usage in visual data analysis [47]. We also include MCA as a DR method designed explicitly for categorical data and, to our knowledge, the only existing technique that does not require a distance matrix but uses the categorical data directly. We create a dissimilarity matrix to compute the projection given one of the described distance measures. The only exception is MCA, as the only dimension reduction method specifically designed for categorical data. The MCA is performed with the set representation of the underlying subset.
@@
Overlap Reduction: Given that some subsets in the categorical data may differ in only one or a few attributes, these subsets will be projected close to each other. This property is desirable in the design of a map by keeping the distances representing similarity coherent. However, it may also introduce overlap if the projected point visually encodes the subset categories through a glyph representation. Additionally, points that are close together will yield small or narrow-shaped Voronoi cells. Thus, we allow for two types of overlap reduction:
Force-directed Graph Drawing: This type of layout applies forces to the nodes and edges of a graph [60]. We add a repulsive force to all points with a strength equal to the radius of the glyph while all points are vertices of a fully connected graph, forcing all points into a configuration without overlap.
Lloyds’s Algorithm: This algorithm relaxes a Voronoi diagram by iteratively moving the point of a cell closer to the centroid of the cell [61], effectively enlarging small cells in each step. We apply iterations of Lloyds’s algorithm until all glyphs can be drawn without overlap.
III-C Representing Categorical Data Subsets in Scatterplots
@@
We implemented the visual components of the Categorical Data Map using D3 [62]. To represent categorical subsets, we developed four glyph representations and the layout enrichment based on experiences gained during the design phase addressing (R3). We use the d3.schemeCategory10 color scale, a well-established color scale for categorical data.
@@
Figure 3: Representation of subsets for a dataset with eight attributes. (a) shows the eight attributes in four equal-sized segments while the size encodes the subset size. (b) shows a similar glyph, but instead, the size is encoded by a bar at the top of the glyph, and all glyphs have the same size. (c) Encodes the attributes similar to the area square but is circle-shaped. (d) encodes the size by an arc filled according to the subset size.
@@
Glyph Representation: To represent categorical subsets, we developed four glyph representations. All glyphs visualize the attributes and their respective values by dividing a square or circle into segments of equal size, such that each segment represents one attribute. This square-based glyph is inspired by pixel visualizations pioneered by Keim et al. [63]. In Fig. 3, this is represented by the categories $a_{1}$ to $a_{8}$ for the case of a dataset with eight attributes. For all glyphs, the segments are colored according to the respective category of the attribute. However, we discuss some limitations in section VI. The area-based glyphs represent the relative size of a subset $s\in\mathbb{N}$ by the area (see Fig. 3 (a) and (c)). Thus, we calculate the width and height accordingly. The bar- and arc-based glyphs have a fixed size to minimize space requirements and overlap issues with neighboring glyphs (see Fig. 3 (b) and (d)). To reduce overlap while preserving the relative proximity of the projected points, we decided to map a subset’s size $s\in\mathbb{N}$ to a bar at the top or an arc sounding the glyph as an alternative encoding for the subset size. Hence, each unique subset is represented by an equally sized square or circle with an indicator filled according to the relative percentage of data points the subset represents. This enables users to perceive similar subsets and assess the size of each group.
@@
Layout Enrichment: To enable the observation of cluster characteristics and explore attributes in the projected space, we show a Voronoi diagram for a selected attribute Fig. 2). The Voronoi diagram automatically partitions the map into polygons such that each polygon contains exactly one subset. By selecting one attribute of interest, the partition for the selected attribute gets displayed in the background of the projection. The color of the polygon then encodes the category of the selected attribute. Thereby, it is possible to directly spot cluster regions for the selected attribute and to identify cluster boundaries and outlying data points. The appearance of the background can differ a lot across attributes (see  Fig. 5). Attributes form distinct contiguous areas of different sizes, indicating a neighborhood or larger area of subsets of the same category. We quantify this property to rank attributes since large contiguous areas are easier to interpret and hint at attributes with a high impact on the projection. We also added detail-on-demand using tooltips, allowing users to see the respective category for each polygon directly.
III-D Interacting with Attributes and Subsets
Our method does not focus on interaction. However, our prototype allows for interactions on attributes and subsets.
Figure 4: Through interaction, the borders of a second attribute can be added to the plot, e.g., Purchaser Living In is shown in the background as the primary attribute and Location of Purchased Property is shown in the foreground.
Attribute Selection: Users can change the attribute visualized through layout enrichment. However, we also show the outline for categories of a second selected attribute (see Fig. 4). We add the borders of categories to the foreground if another attribute is already selected and visualized in the background. This visual cue does allow for the observation of one main attribute and a second attribute, similar to the outline of MosaicSets [64], which introduces less clutter and thus requires less effort to perceive. We initially used textures with different colors to represent different categories. However, using textures of different colors to fill each cell in the Voronoi portioning introduced excessive clutter, and the interpretation of common regions was difficult.
Subset Selection: We allow for the selection and highlighting of groups of subsets. Once the user has selected data items, using Lasso selection, we show the common categories of the selection and highlight all data items outside of the selection with the same combination of categories. This interaction enables cluster analysis since all categories that contribute to the selection are highlighted. Additionally, all subsets matching the common categories of the selection are also highlighted. Together, this allows analysts to observe and judge group cohesion along with the contributing attributes.
III-E Measuring Fracturedness
@@
We quantify fracturedness, generally defined as the strength with which the Voronoi partitioning of an attribute appears disjointed and fractured (see Fig. 5). We use fracturedness to suggest attributes for analysis, e.g., the lower the fracturedness value, the larger the continus areas of categories and thus the more straightforward to orient along addressing (R4). We use the Delaunay triangulation of the Voronoi diagram as a basis for our measures [66]. Before describing the measures, we define the common notations following established notations [67, 68]. Let $G\coloneqq(V,E)$ be the Delaunay triangulation of the discrete set of points $P$ resulting from the projection (see subsection III-B). Thus, $G$ is an undirected graph and the dual graph of the Voronoi diagram of the points $P$ Therefore, there exists exactly one $v\in V$ for every $p\in P$ defining its x,y-location, and categories. Each vertex $v\in V$ has exactly one associated category $\mathscr{C}_{n}(v)\in\mathcal{C}_{n}$ for each attribute $a_{n}\in\mathcal{A}$.
Edge-based Fracturedness: We measure the number of edges in $G$ that connect cells with different associated attributes. This concept is shown Fig. 6. We define an edge $e\in E$ as $\{v_{1},v_{2}\}$ with $v_{1},v_{2}\in V$ and $v_{1}\neq v_{2}$. An edge contributes to fracturedness, if the category for the analyzed attribute $a_{n}$ and its associated categories in $\mathcal{C}_{n}$ differs for the connected vertices, i.e., $\mathscr{C}_{n}(v_{1})\neq\mathscr{C}_{n}(v_{2})$ for $\{v_{1},v_{2}\}\in E$.
$\mathpzc{f}_{\textrm{edge}}(\{v_{1},v_{2}\})\coloneqq\begin{cases}0:\mathscr{C}_{n}(v_{1})=\mathscr{C}_{n}(v_{2})\\ 1:\mathscr{C}_{n}(v_{1})\neq\mathscr{C}_{n}(v_{2})\end{cases}$ (5)
@@
Figure 6: We illustrate edge-based fracturedness with a Delaunay triangulation, shown in black, and a Voronoi partitioning with cell borders shown in red. The cells are colored according to the categories of an attribute. $v_{1}$, $v_{2}$ and $v_{3}$ are vertices of the Delaunay triangulation. The edge ${v_{1},v_{2}}$ will not contribute to edge-based fracturedness, since it connects cells representing the same categories of a given attribute. Edge ${v_{2},v_{3}}$ contributes to edge-based fracturedness, because it connects cells representing different categories.
Component-based Fracturedness: This measure quantifies the number of continuous areas an attribute produces in the plot through its categories. We show the concept of component-based fracturedness in Fig. 7. Each category $c\in\mathcal{C}_{n}$ defines an induced subgraph $G[S(c)]$ of $G$, with $S(c)\subset V$ for all $c\in\mathcal{C}_{n}$ of an attribute $a_{n}\in\mathcal{A}$. The induced subgraph $G[S(c)]$ is a graph with the vertices $S(c)$ and the edges in $E$ with both of its vertices in $S(c)$. We formally define $S(c)$ for a category $c\in\mathcal{C}_{n}$ in Equation 7.
@@
With this definition, a category defines a partition of $V$, i.e., $\bigcup_{c\in\mathcal{C}_{n}}S(c)=V$ and a vertex $v\in V$ can only have one category $\mathscr{C}_{n}(v)$, thus $\bigcap_{c\in\mathcal{C}_{n}}S(c)=\emptyset$ for a given attribute $a_{n}$. Therefore, there exits $\left|\mathcal{C}_{n}\right|$ subgraphs of $G$ for attribute $a_{n}\in\mathcal{A}$. Let $\omega(G)$ be the number of connected components of any graph $G$. The component-based fracturedness is dependent on the number of connected components of all subgraphs $\omega(G[S(c)])$ for each $c\in\mathcal{C}_{n}$. We define the sum of the number of components of all induced subgraphs as $\Omega(a_{n})$ for an attribute $a_{n}\in\mathcal{A}$. $\Omega(a_{n})$ is formally defined in Equation 8.
@@
We can also quantify the fracturedness a single category contributes to the overall measure. This allows us to differentiate categories forming contiguous areas and highly fractured ones. The fracturedness $\mathpzc{f}_{\textrm{comp}}(c)$ of a single category $c\in\mathcal{C}_{n}$ is defined in Equation 9.
@@
The sum of all components-based fracturedness values of individual categories $c\in\mathcal{C}_{n}$ is summing to the fracturedness of the attribute $a_{n}\in\mathcal{A}$ the above value. We express this relationship in Equation 11.
@@
A mathematical proof of the equivalence described in Equation 11 is available in the supplementary material.
Figure 7: We describe component-based fracturedness with a Voronoi partitioning with cell borders shown in red. The associated Delaunay triangulation is shown in black. The cells are colored according to the categories of an attribute. $\omega_{1}$ to $\omega_{5}$ are five components induced by an attribute through the sub-graphs associated with a category. Solid lines connect each sub-graph, while dashed lines are not part of any sub-graph.
IV Interpreting the Categorical Data Map

### v4

@@ 2.1 Projecting Categorical Data
The Categorical Data Map enables the visual clustering of similar categorical subsets and separating outliers, addressing (C1) and (C2). At the core, we rely on DR to create a scatterplot-like visualization. In general, we describe encoding $E$, distance measure $M$, DR method $P$, and overlap reduction method $O$ to project a categorical dataset $x$ by applying $O(P(M(E(x))))$.
@@
Encoding (E): We convert all data items into a set representing their categorical data values. We define the set of all attributes as $\mathcal{A}\coloneqq\{a_{1},a_{2},\dots,a_{\left|\mathcal{A}\right|}\}$ and the possible categories associated with attribute $a_{i}$ as the set $\mathcal{C}_{i}\coloneqq\{c_{i}^{1},c_{i}^{2},\dots,c_{i}^{\left|\mathcal{C}_{i}\right|}\}$ with $i\in\mathbb{N}$. $\left|\mathcal{A}\right|$ is the cardinality of a set representing a data item, i.e., the number of attributes since a data item has one category associated with each attribute. We denote a data item as $x_{n}=(c_{1}^{n_{1}},c_{2}^{n_{2}},...,c_{\left|\mathcal{A}\right|}^{n_{\left|\mathcal{A}\right|}})$. From a practical point of view, we make sure that all categories have a unique descriptor across all attributes. We then create a representation compatible with the distance measure. We explored the set representation and two variants of one-hot encoding [20, 35] (see supplementary material for more details).
Distance Measure (M): With the set representation, we can describe the categories of a data item to define similarity. Based on surveys on distance measures for categorical data [22, 17, 76], we chose and evaluated three set-based distance measures: Overlap coefficient [83], Jaccard Similarity Index [39], and Sørenson-Dice coefficient [73]. By including one-hot encoding, converting each categorical value to a new binary dimension enables us to use classical distance measures, such as Euclidean or Manhattan distance, to describe a dissimilarity relationship (see supplementary material for more details).
@@
Projection Method (P): DR techniques are a set of non-/linear transformation methods with which a dataset’s dimensionality can be reduced. We compared the following two DR methods.
@@
Multiple Correspondence Analysis (MCA): This method is the categorical equivalent of PCA. MCA creates groups of items that are similar according to their categories. Objects sharing the same categories are placed close together, and objects with differing categories are placed far apart [34]. To our knowledge, MCA is the only existing technique that directly uses the set representation of categorical data.
Multidimensional Scaling (MDS): This method describes a set of linear and nonlinear DR techniques that attempt to preserve pairwise distances. Multiple criteria are possible; Kruskal’s stress optimization criterion is usually used [50]. We create a dissimilarity matrix to compute the projection given one of the described distance measures.
We chose the two methods based on their popularity and common usage in visual data analysis [19, 30] and compared MDS and MCA as DR methods for categorical data.
However, a key difference between both methods is that MCA reduces the number of projected points to the number of unique subsets, while MDS, applied naively, would result in a number of projected points equal to the number of categorial data items. Since categorical datasets can contain many duplicates, projecting each data point individually and using a DR method for numeric data (e.g., MDS) could lead to multiple data points being projected to the same position. The main reason is that the distance of identical points is zero. To achieve a comparable result, i.e., the same number of projected points, we remove all duplicates and project one data point for each unique combination of attribute values, i.e., for each categorical subset, describing the prototype of the represented data subset. A second reason is that we want to show the subsets represented by a point irrespective of the method (e.g., MCA or MDS). We visually represent a subset’s size (see Sec. 2.2). Reducing the number of data points also improves the runtime of projection algorithms for datasets with duplicate items.
@@
Overlap Reduction (O): Given that some subsets in the categorical data may differ in only one or a few attributes, these subsets will be projected close to each other. This property is desirable in the design of a map by keeping the distances representing similarity coherent. However, it may also introduce overlap if the projected point visually encodes the subset categories through a glyph representation. Additionally, points that are close together will yield small or narrow-shaped Voronoi cells. Thus, we allow users to reduce the overlap after projecting the data using a method based on force-directed graph drawing. This type of layout applies forces to the nodes and edges of a graph [47]. We add a repulsive force to all points with a strength equal to the radius of the glyph while all points are vertices of a fully connected graph, forcing all points into a configuration without overlap but with minimal space in between the glyphs.
2.2 Representing Categorical Data Subsets in Scatterplots
@@
We implemented the visual components of the Categorical Data Map using D3 [18]. To represent categorical subsets, we developed four glyph representations and the layout enrichment based on experiences gained during the design phase, addressing (C3). To represent categories, we use the d3.schemeCategory10 color scale, a well-established color scale for categorical data.
@@
Figure 1: Representation of subsets for a dataset with eight attributes. (a) shows the eight attributes in four segments with the same area while the size encodes the overall subset size. (b) shows a similar glyph, but instead, the size is encoded by a bar at the top, and all glyphs have the same size. (c) Encodes the attributes similar to the area square but is circle-shaped. (d) encodes the size by an arc filled according to the subset size.
@@
Glyph Representation: To represent categorical subsets, we developed four glyph representations. All glyphs visualize the attributes and their respective values by dividing a square or circle into segments of equal size, such that each segment represents one attribute. This square-based glyph is inspired by pixel visualizations pioneered by Keim et al. [45]. In Fig. 1, this is represented by the categories $a_{1}$ to $a_{8}$ for the case of a dataset with eight attributes. For all glyphs, the segments are colored according to the respective category of the attribute. However, we discuss some limitations in Sec. 5. The area-based glyphs represent the relative size of a subset $s\in\mathbb{N}$ by the area (see Fig. 1 (a) and (c)). Thus, we calculate the width and height accordingly. The bar- and arc-based glyphs have a fixed size to minimize space requirements and overlap issues with neighboring glyphs (see Fig. 1 (b) and (d)). To reduce overlap while preserving the relative proximity of the projected points, we decided to map a subset’s size $s\in\mathbb{N}$ to a bar at the top or an arc surrounding the glyph as an alternative encoding for the subset size. Hence, each unique subset is represented by a square or circle sized relative to the percentage of data points the subset represents or an indicator filled accordingly. This enables users to perceive similar subsets and assess the size of each group.
@@
Layout Enrichment: To enable the observation of cluster characteristics and explore attributes in the projected space, we show a Voronoi diagram [9] for a selected attribute (see The Categorical Data Map: A Multidimensional Scaling-Based Approach). The Voronoi diagram automatically partitions the map into polygons such that each polygon contains exactly one subset. By selecting one attribute of interest, the partition for the selected attribute gets displayed in the background of the projection. The color of the polygon then encodes the category of the selected attribute. Thereby, it is possible to directly spot cluster regions for the selected attribute and to identify cluster boundaries and outlying data points. The appearance of the background can differ a lot across attributes (see  Fig. 2). Attributes form distinct contiguous areas of different sizes, indicating a neighborhood or larger area of subsets of the same category. We added detail-on-demand using tooltips, allowing users to see the respective category for each polygon directly.
2.3 Measuring Fracturedness
@@
We quantify fracturedness, generally defined as the strength with which the Voronoi partitioning of an attribute appears disjointed and fractured (see Fig. 2). We use fracturedness to suggest attributes for analysis, e.g., the lower the fracturedness value, the larger the contiguous areas of categories and thus the more straightforward to orient along, addressing (C4). We use the Delaunay triangulation of the Voronoi diagram [9] as a basis for our measures. In contrast to Aupetit and Catz [7], we describe measures for purely categorical datasets. Before describing the measures, we define the common notations following established notations [24, 7]. Let $G\coloneqq(V,E)$ be the Delaunay triangulation of the discrete set of points $P$ resulting from the projection (see Sec. 2.1). Thus, $G$ is an undirected graph and the dual graph of the Voronoi diagram of the points $P$. Therefore, there exists exactly one $v\in V$ for every $p\in P$ defining its x,y-location and categories. Each vertex $v\in V$ has exactly one associated category $\mathscr{C}_{n}(v)\in\mathcal{C}_{n}$ for each attribute $a_{n}\in\mathcal{A}$.
@@
[…]
Figure 3: We illustrate edge-based fracturedness with a Delaunay triangulation shown in black, and a Voronoi partitioning with cell borders shown in red. The cells are colored according to the categories of an attribute. $v_{1}$, $v_{2}$ and $v_{3}$ are vertices of the Delaunay triangulation. The edge ${v_{1},v_{2}}$ will not contribute to edge-based fracturedness, since it connects cells representing the same category of a given attribute. Edge ${v_{2},v_{3}}$ contributes to edge-based fracturedness because it connects cells representing different categories.
Component-based Fracturedness: This measure quantifies the number of continuous areas an attribute produces in the plot through its categories. We show the concept of component-based fracturedness in Fig. 4. Each category $c\in\mathcal{C}_{n}$ defines an induced subgraph $G[S(c)]$ of $G$, with $S(c)\subset V$ for all $c\in\mathcal{C}_{n}$ of an attribute $a_{n}\in\mathcal{A}$. The induced subgraph $G[S(c)]$ is a graph with the vertices $S(c)$ and the edges in $E$ with both of its vertices in $S(c)$. We formally define $S(c)$ for a category $c\in\mathcal{C}_{n}$ in Eq. 2.
@@
With this definition, a category defines a partition of $V$, i.e., $\bigcup_{c\in\mathcal{C}_{n}}S(c)=V$ and a vertex $v\in V$ can only have one category $\mathscr{C}_{n}(v)$, thus $\bigcap_{c\in\mathcal{C}_{n}}S(c)=\emptyset$ for a given attribute $a_{n}$. Therefore, there exits $\left|\mathcal{C}_{n}\right|$ subgraphs of $G$ for attribute $a_{n}\in\mathcal{A}$. Let $\omega(G)$ be the number of connected components of any graph $G$. The component-based fracturedness is dependent on the number of connected components of all subgraphs $\omega(G[S(c)])$ for each $c\in\mathcal{C}_{n}$ (see $s_{1}$ to $s_{6}$ in Fig. 4). We define the sum of the number of components of all induced subgraphs as $\Omega(a_{n})$ for an attribute $a_{n}\in\mathcal{A}$. $\Omega(a_{n})$ is formally defined in Eq. 3:
@@
We can also quantify the fracturedness a single category contributes to the overall measure. This allows us to differentiate categories forming contiguous areas and highly fractured ones. The fracturedness $\mathpzc{f}_{\textrm{comp}}(c)$ of a single category $c\in\mathcal{C}_{n}$ is defined in Eq. 4:
@@
The sum of all component-based fracturedness values of individual categories $c\in\mathcal{C}_{n}$ is equal to the fracturedness of the attribute $a_{n}\in\mathcal{A}$. We express this relationship in Eq. 6:
@@
A mathematical proof of the equivalence described in Eq. 6 can be found in the supplementary material.
[…]
2.4 Interacting with Attributes and Subsets
Our prototype allows interactions on the attributes of the dataset shown in the side panel and projected subsets.
Attribute Selection: Users can change the attribute visualized through layout enrichment. We also show the outline for categories of a second selected attribute (see Fig. 5). We add the borders of categories to the foreground if another attribute is already selected and visualized in the background. This visual cue does allow for the observation of one main attribute and a second attribute, similar to the outline of MosaicSets [69]. This introduces less clutter and thus requires less effort to perceive. We initially used textures with different colors to represent different categories. However, using textures of different colors to fill each cell in the Voronoi portioning introduced excessive clutter, and the interpretation of common regions was difficult.
Subset Selection: We allow for the selection and highlighting of groups of subsets. Once the user has selected data items, we show the common categories of the selection using Lasso selection and highlight all data items outside of the selection with the same combination of categories in the side panel on the left, similar to the proximity visualization for continuous data proposed by Aupetit and Catz [6]. This interaction enables cluster analysis since all common categories among the selected items are highlighted (see side panel in Fig. 5). Thus, visual groupings can be compared with respect to the categories and attributes contributing to cluster cohesion. Additionally, all subsets matching the common categories of the selection are also highlighted (see plot in Fig. 5). Together, this allows analysts to observe and judge group cohesion along with the contributing attributes.
Attribute and Category Ordering: A user can select attributes of the dataset listed on the side panel to change the attribute encoded in the foreground and background of the plot. By default, attributes are sorted by their edge-based fracturedness in ascending order, and categories are ordered by their individual contributions to component-based fracturedness in ascending order, allowing for a focus on attributes forming clear splits in the projection space. When selecting subsets (see previous paragraph), the lists of common attributes and distinct attributes are also ordered similarly.
Figure 5: Through user selection, the borders of a second attribute can be added to the foreground of the plot, e.g., Purchaser Currently Living In is shown in the background as the primary attribute, and Location of Purchased Property is shown in the foreground.
3 Interpreting the Categorical Data Map